OpenStudy (anonymous):

Limit Help! Rationalize the numerator to find the exact value of the limit analytically: f(x)= the square root of (2x+1) minus the square root of 3 all of that over (x-1)

OpenStudy (anonymous):

\[\frac{\sqrt{2x+1}-3}{x-1}\]?

OpenStudy (anonymous):

root (2x+1) - root 3

OpenStudy (anonymous):

do you know what it means when it says "rationalize the numerator"? it means multiply top and bottom by \(\sqrt{2x+1}+\sqrt3\) leave the denominator in factored form

OpenStudy (anonymous):

the numerator will be \[2x+1-3=2x-2=2(x-1)\]

OpenStudy (anonymous):

the denominator will be \[(x-1)(\sqrt{2x+1}+\sqrt3)\]

OpenStudy (anonymous):

i got that much, but I didn't know how to solve the denominator

OpenStudy (anonymous):

leave it be

OpenStudy (anonymous):

you get \[\frac{2(x-1)}{(x-1)(\sqrt{2x+1}+3)}\]

OpenStudy (anonymous):

then cancel and get \[\frac{2}{\sqrt{2x+1}+\sqrt3}\]

OpenStudy (anonymous):

then replace \(x\) by \(1\) and you are done

OpenStudy (anonymous):

why do you replace x with 1?

OpenStudy (anonymous):

didn't you want the limit as \(x\to 1\)?

OpenStudy (anonymous):

it didn't give me what x was approaching - it just said "Rationalize the numerator to find the exact value of the limit analytically"

OpenStudy (anonymous):

the answer in the back of the book says the limit is root 3 divided by 3

OpenStudy (anonymous):

up until this point \[\frac{2(x-1)}{(x-1)(\sqrt{2x+1}+3)}\] you cannot replace \(x\) by \(1\) because you will get \(\frac{0}{0}\) but once you get to \[\frac{2}{\sqrt{2x+1}+\sqrt3}\] you can you get just what it has in the back of the book

OpenStudy (anonymous):

it doesn't give me what to replace x with is what I'm saying

OpenStudy (anonymous):

oh wait. I'm sorry - I didn't see the top of the page. x approaches 1+

OpenStudy (anonymous):

@satellite73